OpenStudy (anonymous):

Nathaniel can weld a railing in 75 minutes. Brenda can weld a railing 25 minutes faster. If they work together, how many minutes does it take them to weld the railing?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

your idea?

OpenStudy (anonymous):

I would do like 1/75+1/25=1/x

OpenStudy (anonymous):

bcuz 100 min. isn't an answer choice lol

OpenStudy (loser66):

For Nathaniel, he needs 75 min to finish the job. That is in 1 minute, he finishes 1/75 part of the job, right? Same as Brenda with 1/25 so that in 1 minute, both them will finish \(\dfrac{1}{75}+\dfrac{1}{25}= \dfrac{4}{75}\) part of the job. Summary: 1min------------4/75 ? min-----------75/75??

OpenStudy (anonymous):

50 ??

OpenStudy (loser66):

nope 75/4 = 72 minutes + 3/4 minutes. That is 72minutes 45seconds

OpenStudy (anonymous):

25 minutes 30 minutes 50 minutes 150 minutes

OpenStudy (anonymous):

Those are the answers. I have no idea what to do

OpenStudy (loser66):

oh!!@ sorry 75/4 = 18 + 3/4

OpenStudy (anonymous):

:) It's okay hahha

OpenStudy (loser66):

But it is just 18 minutes 45seconds which is not one of the options. ha!!

OpenStudy (anonymous):

lolll I thought so too! So I guess it would be 25 min.???

OpenStudy (loser66):

I don't guess but don't know what is wrong with the logic. Ok, let's get help @ganeshie8 Please, explain me what is wrong.!!

OpenStudy (anonymous):

Lol thank-you :D

ganeshie8 (ganeshie8):

`Nathaniel can weld a railing in 75 minutes. Brenda can weld a railing 25 minutes faster.` That means Brenda can finish the work in `50` minutes right ?

ganeshie8 (ganeshie8):

so we must be solving \[\dfrac{1}{75}+\dfrac{1}{50} = \dfrac{1}{x}\]

OpenStudy (loser66):

YYYYYYYYYYYES

OpenStudy (anonymous):

:) I agree. So would it be 50 min.?

OpenStudy (anonymous):

Guys?

ganeshie8 (ganeshie8):

what is 50 ?

OpenStudy (anonymous):

No idea. Thanks @ganeshie8 and @Loser66 :) !!!!!

OpenStudy (loser66):

hehehehe..... no idea??? That is we are cheated by you??